# Sheffer polynomials and the s-ordering of exponential boson operators

### v1

@@ Abstract
The $s$-ordered form of any product of single-mode boson creation and annihilation operators, containing only a single annihilator, is computed explicitly. The $s$-ordering concept originated in quantum optics, and subsumes normal, symmetric (Weyl), and anti-normal ordering for any two operators satisfying a canonical commutation relation. Because the $s$-ordering map can be viewed as producing a function of a complex variable, its inverse is a quantization map that takes such “classical” functions to quantum operators. The explicit $s$-ordered expressions are derived with the aid of a parametric family of Sheffer polynomial sequences (or equivalently an exponential Riordan array of polynomial coefficients), called the Hsu–Shiue family. To yield orderings interpolating between normal and anti-normal, this family is extended in an intricate way.
@@ 1 Introduction
${\rm e}^{\lambda({a^{\dagger}}a)}=\left\{\begin{aligned} &\mathopen{:}\,\mathrel{\exp\left[\left({\rm e}^{\lambda}-1\right){a^{\dagger}}a\right]}\,\mathclose{:}_{N},\\ &\mathopen{:}\,\mathrel{\frac{2}{1+{\rm e}^{\lambda}}\exp\left[\frac{2({\rm e}^{\lambda}-1)}{1+{\rm e}^{\lambda}}{a^{\dagger}}a\right]}\,\mathclose{:}_{W},\\ &\mathopen{:}\,\mathrel{{\rm e}^{-\lambda}\exp\left[\left(1-{\rm e}^{\lambda}\right){a^{\dagger}}a\right]}\,\mathclose{:}_{A}.\end{aligned}\right$. (1)
@@
${\rm e}^{\lambda({a^{\dagger}}a)}=\mathopen{:}\,\mathrel{\frac{2}{1+{\rm e}^{\lambda}-s(1-{\rm e}^{\lambda})}\exp\left[\frac{2({\rm e}^{\lambda}-1)}{1+{\rm e}^{\lambda}-s(1-{\rm e}^{\lambda})}{a^{\dagger}}a\right]}\,\mathclose{:}_{s},$ (2)
which subsumes the preceding three.
@@
This $F^{(A)}$, which is obtained (by setting $s=\nobreak+1$) from a general $s$-parametrized representation of $\hat{\rho}$, which is denoted by $F^{(s)}$, may not be non-negative on the $\alpha$-plane; in which case it cannot be interpreted as a classical probability density, and the state is said to be non-classical. But non-classicality can be quantified: as $s$ is lowered from +1+1 toward 0 and 1-1, regions of negativity (if any) of $F^{(s)}$ and its marginals shrink at some rate and vanish [6].
Quantum optics models in which the time evolution of $s$-parametrized quasiprobability densities has been studied include the widely applicable Jaynes–Cummings model of a two-level atom (a qubit) interacting with a single light mode of a cavity [7], and the Drummond–Walls model of dispersive optical bistability [8]. Plots of evolving densities, at multiple alternative values of $s$, reveal much about model dynamics.
In this paper, we go beyond previous work on the foundations of $s$-ordering by finding closed-form, analytic expressions of $s$-ordered representations denoted by $F^{(s)}$ (i.e., $s$-dependent functions on the complex $\alpha$-plane), for certain single-mode exponential boson operators of the form ${\rm e}^{\lambda\omega}$, where $\omega\in\mathcal{W}$. That is, for each $s$ we find an $F^{(s)}$ with $s$-quantization equal to $\omega$.
The case most amenable to an analytic treatment is when $\omega$ is a monomial or “boson string”: a product of $a$’s and $a^{\dagger}$}’s, i.e., a word over $\{a,{a^{\dagger}}\}$. The “single annihilator” word $\omega={a^{\dagger}}^{L}a{a^{\dagger}}^{R}$ will be treated, where $e=L+\nobreak R-\nobreak 1\geqslant 0$ is the excess of $a^{\dagger}$}’s over $a$’s. The Cahill–Glauber result, eq. (2) above, is the $s$-representation of ${\rm e}^{\lambda\omega}$ when $(L,R)=(1,0)$ and $\omega={a^{\dagger}}a$.
The chief tool is the theory of Sheffer polynomials [9, 10], or equivalently the theory of exponential Riordan arrays (infinite triangular arrays of polynomial coefficients) [11, 12, 13]. A Sheffer polynomial sequence is a sequence of polynomials $s_{n}(t)$ with $s_{0}(t)\equiv 1$ and $\deg s_{n}=n$, $n\geqslant 0$, which can be constructed in any of several ways, including the use of differential ladder operators that increment and decrement $n$. By extending previous work of Lang [14, 15] and Błasiak and collaborators [16, 17], we first derive normal and anti-normal representations that subsume the first and third identities in (1), namely
$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}&=\mathopen{:}\,\mathrel{\mathcal{HS}({a^{\dagger}}a,\lambda{a^{\dagger}}^{e};\,-e,1,R)}\,\mathclose{:}_{N}\\ &=\mathopen{:}\,\mathrel{(1-e\lambda{a^{\dagger}}^{e})^{-R/e}\exp\left\{\left[\bigl{(}1-e\lambda{a^{\dagger}}^{e}\bigr{)}^{-1/e}-1\right]{a^{\dagger}}a\right\}}\,\mathclose{:}_{N}\end{aligned}$ (3)
@@
$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}&=\mathopen{:}\,\mathrel{\mathcal{HS}(a{a^{\dagger}},\lambda{a^{\dagger}}^{e};\,e,-1,-L)}\,\mathclose{:}_{A}\\ &=\mathopen{:}\,\mathrel{\exp\left\{-a{a^{\dagger}}\left[\bigl{(}1+e\lambda{a^{\dagger}}^{e}\bigr{)}^{-1/e}-1\right]\right\}(1+e\lambda{a^{\dagger}}^{e})^{-L/e}}\,\mathclose{:}_{A},\end{aligned}$ (4)
@@
Our main result (Theorem 5.3) is the explicit $s$-ordering of ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$, valid for arbitrary $s\in\mathbb{C}$. It subsumes eqs. (3) and (4), which are its $s=\nobreak-1$ and $s=\nobreak+1$ special cases, and includes a new and explicit Weyl ordering formula ($s=\nobreak 0$); for the latter, see Theorem 5.5. The general $s$-ordered representation is expressed in terms of what will be called a two-point Hsu–Shiue EGF, interpolating in an intricate sense between the EGF’s in (3) and (4).
@@
The paper is structured as follows. In Sec. 2, operator orderings including $s$-ordering are defined, in part using quantum optics language. Sheffer polynomial sequences and exponential Riordan arrays are introduced in Sec. 3, and the parametric Hsu–Shiue family is defined. Theorem 4.4 subsumes eqs. (3) and (4) above, and serves as a lemma in the derivation by interpolation of the main result, in Sec. 5. Examples are given in Sec. 6.
@@ 2.1 Definitions
The representation as polynomials in $a,{a^{\dagger}}$ is not unique, as the product on $\mathcal{W}$ satisfies $a,{a^{\dagger}}]=1$. $\mathcal{W}$ can be identified with $\mathbb{C}\langle a,{a^{\dagger}}\rangle/\allowbreak(a{a^{\dagger}}-\nobreak{a^{\dagger}}a-\nobreak 1)$, the quotient of the free algebra $\mathbb{C}\langle a,{a^{\dagger}}\rangle$ by the two-sided ideal generated by $a{a^{\dagger}}-\nobreak\nobreak{a^{\dagger}}a-\nobreak 1$. Thus each element of $\mathcal{W}$ is an equivalence class of polynomials in $a,{a^{\dagger}}$, though a distinguished representative can be singled out by imposing an ordering criterion.
@@ 2.2 Orderings
Let $G\in\mathcal{W}$, i.e., let $G(a,{a^{\dagger}})$ be a polynomial in the noncommuting $a,{a^{\dagger}}$ that represents $G$ (and is not unique). One can define two c-number or classical representations (sometimes called transforms) associated to $G$, denoted by $F^{(N)},F^{(A)}\in\mathcal{W}_{c}=\mathbb{C}[\alpha,\alpha^{*}]$, the ring of polynomials in commuting indeterminates $\alpha,\alpha^{*}$, which is $\mathbb{C}\langle\alpha,\alpha^{*}\rangle/\allowbreak(\alpha\alpha^{*}-\nobreak\alpha^{*}\alpha)$. They are defined by requiring that
$G(a,{a^{\dagger}})$ $=\mathopen{:}\,\mathrel{F^{(N)}(\alpha,\alpha^{*})}\,\mathclose{:}_{N},$ (5a)
$G(a,{a^{\dagger}})$ $=\mathopen{:}\,\mathrel{F^{(A)}(\alpha,\alpha^{*})}\,\mathclose{:}_{A},$ (5b)
@@
In the same way, one defines a standardized Weyl-ordered representation $F^{(W)}\in\mathcal{W}_{c}$ of any $G\in\mathcal{W}$ by requiring that $G(a,{a^{\dagger}})$ equal $\mathopen{:}\,\mathrel{F^{(W)}(\alpha,\alpha^{*})}\,\mathclose{:}_{W}$. Here, the Weyl quantization $\mathopen{:}\,\mathrel{\dots}\,\mathclose{:}_{W}$ maps any monomial ${\alpha^{*}}^{n}\alpha^{m}$ to the average of the $\binom{n+m}{n}=\binom{n+m}{m}$ distinct orderings of a product of $n$ $a^{\dagger}$}’s and $m$ $a$’s. (An equivalent definition is that in $\mathcal{W}$, $\mathopen{:}\,\mathrel{(\mu\alpha+\nobreak\nu\alpha^{*})^{p}}\,\mathclose{:}_{W}$ must equal $(\mu a+\nobreak\nu{a^{\dagger}})^{p}$.) For instance,
$\mathopen{:}\,\mathrel{{\alpha^{*}}^{2}\alpha}\,\mathclose{:}_{W}=({a^{\dagger}}^{2}a+{a^{\dagger}}a{a^{\dagger}}+a{a^{\dagger}}^{2})/3,$ (6)
where the right-hand side, an element of $\mathcal{W}$, can be rewritten with the aid of $a,{a^{\dagger}}]=1$ in various ways, including ${a^{\dagger}}^{2}a+\nobreak{a^{\dagger}}$, ${a^{\dagger}}a{a^{\dagger}}$, $a{a^{\dagger}}^{2}-\nobreak{a^{\dagger}}$. So
$G(a,{a^{\dagger}})={a^{\dagger}}a{a^{\dagger}}\quad\Longrightarrow\quad\left\{\begin{aligned} F^{(N)}(\alpha,\alpha^{*})&={\alpha^{*}}^{2}\alpha+\alpha^{*},\\ F^{(W)}(\alpha,\alpha^{*})&={\alpha^{*}}^{2}\alpha,\\ F^{(A)}(\alpha,\alpha^{*})&={\alpha^{*}}^{2}\alpha-\alpha^{*},\end{aligned}\right$. (7)
@@
$D(\beta,\beta^{*})$ $=\exp(\beta{a^{\dagger}}-\beta^{*}a)=\sum_{p=0}^{\infty}\frac{1}{p!}(\beta{a^{\dagger}}-\beta^{*}a)^{p}$ (8)
@@
$\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{W}=\frac{\partial^{n+m}}{\partial\beta^{n}\partial(-\beta^{*})^{m}}D(\beta,\beta^{*})\Bigm{|}_{(\beta,\beta^{*})=(0,0)}.$ (9)
@@
Define an $s$-parametrized deformation of the exponential operator $D(\beta,\beta^{*})$, namely $D(\beta,\beta^{*};s)={\rm e}^{s\beta(-\beta^{*})/2})D(\beta,\beta^{*})$, where $s\in\mathbb{C}$, and for each $n,m$, define the $s$-quantization $\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s}$ of ${\alpha^{*}}^{n}\alpha^{m}$ by
$D(\beta,\beta^{*};s)=\sum_{n,m=0}^{\infty}\frac{\beta^{n}(-{\beta^{*}})^{m}}{n!\,m!}\,\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s},$ (10)
@@
$\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s}=\frac{\partial^{n+m}}{\partial\beta^{n}\partial(-\beta^{*})^{m}}D(\beta,\beta^{*};s)\Bigm{|}_{(\beta,\beta^{*})=(0,0)}.$ (11)
@@
$D(\beta,\beta^{*})=\exp(\beta{a^{\dagger}}-\beta^{*}a)={\rm e}^{s|\beta|^{2}}\mathopen{:}\,\mathrel{{\rm e}^{\beta\alpha^{*}-\beta^{*}\alpha}}\,\mathclose{:}_{s},$ (12)
@@
As an example of (11), one finds with the aid of $a,{a^{\dagger}}]=1$ that
$\mathopen{:}\,\mathrel{{\alpha^{*}}^{2}\alpha}\,\mathclose{:}_{s}={a^{\dagger}}a{a^{\dagger}}+s\,{a^{\dagger}}.$ (13)
@@
$G(a,{a^{\dagger}})=\mathopen{:}\,\mathrel{F^{(s)}(\alpha,\alpha^{*})}\,\mathclose{:}_{s}.$ (14)
@@
$G(a,{a^{\dagger}})={a^{\dagger}}a{a^{\dagger}}\quad\Longrightarrow\quad F^{(s)}(\alpha,\alpha^{*})={\alpha^{*}}^{2}\alpha-s\,\alpha^{*}.$ (15)
When $s=\nobreak 0$, $F^{(s)}$ reduces to the Weyl-ordered representation $F^{(W)}$. It is also not difficult to show, using the Baker–Campbell–Hausdorff formula
${\rm e}^{A}{\rm e}^{B}={\rm e}^{A+B+[A,B]/2},$ (16)
@@
which holds as an equality between power series if $[A,B]$ commutes with $A$ and $B$, that when $s=\nobreak-1$ and $s=\nobreak+1$, $F^{(s)}$ reduces respectively to the normal and anti-normal ordered representations, $F^{(N)}$ and $F^{(A)}$. The s=1,0,+1s=-1,0,+1 cases of (15) are reflected in eq. (7).
@@
In modern quantum optics, the most commonly encountered c-number representations of any element of the Heisenberg–Weyl algebra are the classical functions $F^{(-1)}=F^{(N)}$, $F^{(0)}=F^{(W)}$, $F^{(+1)}=F^{(A)}$ on $\mathbb{C}$ that arise from $G=\hat{\rho}$, where $\hat{\rho}$ is a specified (single-mode) density operator that is an element of $\mathcal{W}$, or an extension of $\mathcal{W}$ that includes formal power series in $a,{a^{\dagger}}$ [2, 21, 22]. These are called the $Q$, $W$, and $P$ quasiprobability densities, and are attributed respectively to Husimi, Wigner, and Glauber–Sudarshan. But in older works on phase-space methods in quantum optics, operators other than density operators and observables were frequently “classicized,” and it was in that general context that $s$-ordering was introduced. (See [23, 24, 25], and more recently [26]. [^1] There are two conventions on the sign of $s$. In the original paper of Cahill–Glauber [1] and a few more recent ones [26, 27, 28], s=+1,1s=+1,-1 correspond to normal and anti-normal ordering; e.g., for a density operator, to $Q$ and $P$. But usually, they correspond to anti-normal and normal ordering (i.e., to $P$ and $Q$). See, e.g., [2, Chapter 12], [6, 29], and [3, Chapter 3]. The latter convention has been adopted here, so the sign of $s$ in eq. (2), taken from [1], has been changed. Note also that in some works (e.g., [21] and [24, Table V]), the meanings of the so-called reciprocal orderings $F^{(N)}$ and $F^{(A)}$, related by negation of $s$, are interchanged.)
@@
Differentiating the exponential operator $D(\beta,\beta^{*};s)={\rm e}^{(s^{\prime}-s)|\beta|^{2}/2}D(\beta,\beta^{*};s^{\prime})$ with respect to $\beta,\beta^{*}$, in a procedure that can be traced to Wilcox [30], and using (11), yields the $s$-ordering conversion formula
@@
$\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s}=\sum_{k=0}^{\min(m,n)}k!\,\binom{n}{k}\binom{m}{k}\left(\frac{s-s^{\prime}}{2}\right)^{k}\mathopen{:}\,\mathrel{{\alpha^{*}}^{n-k}\alpha^{m-k}}\,\mathclose{:}_{s^{\prime}}.$ (17)
The simplest case of (17) is the contraction
$\mathopen{:}\,\mathrel{\alpha^{*}\alpha}\,\mathclose{:}_{s}-\mathopen{:}\,\mathrel{\alpha^{*}\alpha}\,\mathclose{:}_{s^{\prime}}=\frac{s-s^{\prime}}{2},$ (18)
@@
from which the general formula (17) can be derived, much as Wick’s theorem is derived [28]. As another example, the $s=\nobreak+1$, $s^{\prime}=-1$ case of (17) is
$a^{m}{a^{\dagger}}^{n}=\sum_{k=0}^{\min(m,n)}k!\,\binom{n}{k}\binom{m}{k}{a^{\dagger}}^{n-k}a^{m-k},$ (19)
@@
an identity in $\mathcal{W}$ that converts an anti-normal ordering ($s=\nobreak+1$) to a normal one ($s=\nobreak-1$). Thus one can convert between the coefficients of an anti-normal and a normal series, $\sum_{m,n}f_{m,n}^{(A)}a^{m}{a^{\dagger}}^{n}$ and $\sum_{m,n}f_{n,m}^{(N)}{a^{\dagger}}^{n}a^{m}$ [31].

### v2

@@ Abstract
The $s$-ordered form of any product of single-mode boson creation and annihilation operators, containing only a single annihilator, is computed explicitly. The $s$-ordering concept originated in quantum optics, but subsumes normal, symmetric (Weyl), and anti-normal ordering for any two operators satisfying a canonical commutation relation. Because the $s$-ordering map can be viewed as producing a function of a complex variable, its inverse is a quantization map that takes such “classical” functions to quantum operators. The explicit $s$-ordered expressions are derived with the aid of a parametric family of Sheffer polynomial sequences (or equivalently a parametric exponential Riordan array of polynomial coefficients), called the Hsu–Shiue family. To yield orderings interpolating between normal and anti-normal, this family must be extended.
@@ 1 Introduction
${\rm e}^{\lambda({a^{\dagger}}a)}=\left\{\begin{aligned} &\mathopen{:}\,\mathrel{\exp\left[\left({\rm e}^{\lambda}-1\right){a^{\dagger}}a\right]}\,\mathclose{:}_{N},\\ &\mathopen{:}\,\mathrel{\frac{2}{1+{\rm e}^{\lambda}}\exp\left[\frac{2({\rm e}^{\lambda}-1)}{1+{\rm e}^{\lambda}}{a^{\dagger}}a\right]}\,\mathclose{:}_{W},\\ &\mathopen{:}\,\mathrel{{\rm e}^{-\lambda}\exp\left[\left(1-{\rm e}^{-\lambda}\right){a^{\dagger}}a\right]}\,\mathclose{:}_{A}.\end{aligned}\right$. (1.1)
@@
${\rm e}^{\lambda({a^{\dagger}}a)}=\mathopen{:}\,\mathrel{\frac{2}{1+{\rm e}^{\lambda}-s(1-{\rm e}^{\lambda})}\exp\left[\frac{2({\rm e}^{\lambda}-1)}{1+{\rm e}^{\lambda}-s(1-{\rm e}^{\lambda})}{a^{\dagger}}a\right]}\,\mathclose{:}_{s},$ (1.2)
which subsumes the preceding three, and has the quantization
$\mathopen{:}\,\mathrel{{\rm e}^{\lambda({a^{\dagger}}a)}}\,\mathclose{:}_{s}=\frac{2}{2-\lambda-s\lambda}\left(\frac{2+\lambda-s\lambda}{2-\lambda-s\lambda}\right)^{{a^{\dagger}}a}\!=\left[1+\tfrac{1}{2}(1-s)\lambda\right]^{{a^{\dagger}}a}\left[1-\tfrac{1}{2}(1+s)\lambda\right]^{-a{a^{\dagger}}}$ (1.3)
as its inverse.
@@
This $F^{(A)}$, which is obtained (by setting $s=\penalty 10000+1$) from a general $s$-parametrized representation of $\hat{\rho}$, which is denoted by $F^{(s)}$, may not be non-negative on the $\alpha$-plane; in which case it cannot be interpreted as a classical probability density, and the state is said to be non-classical. But non-classicality can be quantified: as $s$ is lowered from +1+1 toward 0 and 1-1, regions of negativity (if any) of $F^{(s)}$ and its marginals shrink at some rate and vanish [6].
Quantum optics models in which the time evolution of $s$-parametrized quasiprobability densities has been studied include the widely applicable Jaynes–Cummings model of a two-level atom (a qubit) interacting with a single light mode of a cavity [7], and the Drummond–Walls model of dispersive optical bistability [8]. Plots of densities evolving with time, at multiple alternative values of $s$, reveal much about model dynamics.
In this paper, we go beyond previous work on the foundations of $s$-ordering by finding closed-form, analytic expressions of $s$-ordered representations denoted by $F^{(s)}$ (i.e., $s$-dependent functions on the complex $\alpha$-plane), for certain single-mode exponential boson operators of the form ${\rm e}^{\lambda\omega}$, where $\omega\in\mathcal{W}$. That is, for each $s$ we find an $F^{(s)}$ for which the $s$-quantization $\mathopen{:}\,\mathrel{F^{(s)}}\,\mathclose{:}_{s}$ equals ${\rm e}^{\lambda\omega}$.
[…]
The chief tool is the theory of Sheffer polynomials [9, 10], or equivalently the theory of exponential Riordan arrays (infinite triangular arrays of polynomial coefficients) [11, 12, 13]. A Sheffer polynomial sequence is a sequence of polynomials $s_{n}(t)$ with $s_{0}(t)\equiv 1$ and $\deg s_{n}=n$, $n\geqslant 0$, which can be constructed in any of several ways, including the use of differential ladder operators that increment and decrement $n$. By extending previous work of Lang [14, 15] and Błasiak and collaborators [16, 17], we first derive normal and anti-normal representations that subsume the first and third identities in (1.1), namely
$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}&=\mathopen{:}\,\mathrel{\mathcal{HS}({a^{\dagger}}a,\lambda{a^{\dagger}}^{e};\,-e,1,R)}\,\mathclose{:}_{N}\\ &=\mathopen{:}\,\mathrel{(1-e\lambda{a^{\dagger}}^{e})^{-R/e}\exp\left\{\left[\bigl(1-e\lambda{a^{\dagger}}^{e}\bigr)^{-1/e}-1\right]{a^{\dagger}}a\right\}}\,\mathclose{:}_{N}\end{aligned}$ (1.4)
@@
$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}&=\mathopen{:}\,\mathrel{\mathcal{HS}(a{a^{\dagger}},\lambda{a^{\dagger}}^{e};\,e,-1,-L)}\,\mathclose{:}_{A}\\ &=\mathopen{:}\,\mathrel{\exp\left\{-a{a^{\dagger}}\left[\bigl(1+e\lambda{a^{\dagger}}^{e}\bigr)^{-1/e}-1\right]\right\}(1+e\lambda{a^{\dagger}}^{e})^{-L/e}}\,\mathclose{:}_{A},\end{aligned}$ (1.5)
@@
[…]
@@
The paper is structured as follows. In Sec. 2, operator orderings including $s$-ordering are defined, in part using quantum optics language. Sheffer polynomial sequences and exponential Riordan arrays are introduced in Sec. 3, and the parametric Hsu–Shiue family is defined. Theorem 4.4 subsumes eqs. (1.4) and (1.5) above, and serves as a lemma in the derivation by interpolation of the main result, in Sec. 5. Examples are given in Sec. 6.
@@ 2.1 Definitions
[…]
@@ 2.2 Orderings
Let $G\in\mathcal{W}$, i.e., let $G(a,{a^{\dagger}})$ be a polynomial in the noncommuting $a,{a^{\dagger}}$ that represents $G$ (and is not unique). One can define two c-number or classical representations (sometimes called transforms) associated to $G$, denoted by $F^{(N)},F^{(A)}\in\mathcal{W}_{c}=\mathbb{C}[\alpha,\alpha^{*}]$, the ring of polynomials in commuting indeterminates $\alpha,\alpha^{*}$, which is $\mathbb{C}\langle\alpha,\alpha^{*}\rangle/\allowbreak(\alpha\alpha^{*}-\penalty 10000\alpha^{*}\alpha)$. They are defined by requiring that
$G(a,{a^{\dagger}})$ $=\mathopen{:}\,\mathrel{F^{(N)}(\alpha,\alpha^{*})}\,\mathclose{:}_{N},$ (2.1a)
$G(a,{a^{\dagger}})$ $=\mathopen{:}\,\mathrel{F^{(A)}(\alpha,\alpha^{*})}\,\mathclose{:}_{A},$ (2.1b)
@@
In the same way, one defines a standardized Weyl-ordered representation $F^{(W)}\in\mathcal{W}_{c}$ of any $G\in\mathcal{W}$ by requiring that $G(a,{a^{\dagger}})$ equal $\mathopen{:}\,\mathrel{F^{(W)}(\alpha,\alpha^{*})}\,\mathclose{:}_{W}$. Here, the Weyl quantization $\mathopen{:}\,\mathrel{\dots}\,\mathclose{:}_{W}$ maps any monomial ${\alpha^{*}}^{n}\alpha^{m}$ to the average of the $\binom{n+m}{n}=\binom{n+m}{m}$ distinct orderings of a product of $n$ $a^{\dagger}$}’s and $m$ $a$’s. (An equivalent definition is that in $\mathcal{W}$, $\mathopen{:}\,\mathrel{(\mu\alpha+\penalty 10000\nu\alpha^{*})^{p}}\,\mathclose{:}_{W}$ must equal $(\mu a+\penalty 10000\nu{a^{\dagger}})^{p}$.) For instance,
$\mathopen{:}\,\mathrel{{\alpha^{*}}^{2}\alpha}\,\mathclose{:}_{W}=({a^{\dagger}}^{2}a+{a^{\dagger}}a{a^{\dagger}}+a{a^{\dagger}}^{2})/3,$ (2.2)
where the right-hand side, an element of $\mathcal{W}$, can be rewritten with the aid of $a,{a^{\dagger}}]=1$ in various ways, including ${a^{\dagger}}^{2}a+\penalty 10000{a^{\dagger}}$, ${a^{\dagger}}a{a^{\dagger}}$, $a{a^{\dagger}}^{2}-\penalty 10000{a^{\dagger}}$. So
$G(a,{a^{\dagger}})={a^{\dagger}}a{a^{\dagger}}\quad\Longrightarrow\quad\left\{\begin{aligned} F^{(N)}(\alpha,\alpha^{*})&={\alpha^{*}}^{2}\alpha+\alpha^{*},\\ F^{(W)}(\alpha,\alpha^{*})&={\alpha^{*}}^{2}\alpha,\\ F^{(A)}(\alpha,\alpha^{*})&={\alpha^{*}}^{2}\alpha-\alpha^{*},\end{aligned}\right$. (2.3)
@@
$D(\beta,\beta^{*})$ $=\exp(\beta{a^{\dagger}}-\beta^{*}a)=\sum_{p=0}^{\infty}\frac{1}{p!}(\beta{a^{\dagger}}-\beta^{*}a)^{p}$ (2.4)
@@
$\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{W}=\frac{\partial^{n+m}}{\partial\beta^{n}\partial(-\beta^{*})^{m}}D(\beta,\beta^{*})\Bigm|_{(\beta,\beta^{*})=(0,0)}.$ (2.5)
@@
Define an $s$-parametrized deformation of the exponential operator $D(\beta,\beta^{*})$, namely $D(\beta,\beta^{*};s)={\rm e}^{s\beta(-\beta^{*})/2}D(\beta,\beta^{*})$, where $s\in\mathbb{C}$; and for each $n,m$, define the $s$-quantization $\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s}$ of ${\alpha^{*}}^{n}\alpha^{m}$ by
$D(\beta,\beta^{*};s)=\sum_{n,m=0}^{\infty}\frac{\beta^{n}(-{\beta^{*}})^{m}}{n!\,m!}\,\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s},$ (2.6)
@@
$\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s}=\frac{\partial^{n+m}}{\partial\beta^{n}\partial(-\beta^{*})^{m}}D(\beta,\beta^{*};s)\Bigm|_{(\beta,\beta^{*})=(0,0)}.$ (2.7)
@@
$D(\beta,\beta^{*})=\exp(\beta{a^{\dagger}}-\beta^{*}a)={\rm e}^{s|\beta|^{2}/2}\mathopen{:}\,\mathrel{{\rm e}^{\beta\alpha^{*}-\beta^{*}\alpha}}\,\mathclose{:}_{s},$ (2.8)
@@
As an example of (2.7), one finds with the aid of $a,{a^{\dagger}}]=1$ that
$\mathopen{:}\,\mathrel{{\alpha^{*}}^{2}\alpha}\,\mathclose{:}_{s}={a^{\dagger}}a{a^{\dagger}}+s\,{a^{\dagger}}.$ (2.9)
@@
$G(a,{a^{\dagger}})=\mathopen{:}\,\mathrel{F^{(s)}(\alpha,\alpha^{*})}\,\mathclose{:}_{s}.$ (2.10)
@@
$G(a,{a^{\dagger}})={a^{\dagger}}a{a^{\dagger}}\quad\Longrightarrow\quad F^{(s)}(\alpha,\alpha^{*})={\alpha^{*}}^{2}\alpha-s\,\alpha^{*}.$ (2.11)
[…]
${\rm e}^{A}{\rm e}^{B}={\rm e}^{A+B+[A,B]/2},$ (2.12)
@@
which holds as an equality between power series if $[A,B]$ commutes with $A$ and $B$, that when $s=\penalty 10000-1$ and $s=\penalty 10000+1$, $F^{(s)}$ reduces respectively to the normal and anti-normal ordered representations, $F^{(N)}$ and $F^{(A)}$. The s=1,0,+1s=-1,0,+1 cases of (2.11) are reflected in eq. (2.3).
@@
In modern quantum optics, the most commonly encountered c-number representations of any element of the Heisenberg–Weyl algebra are the classical functions $F^{(-1)}=F^{(N)}$, $F^{(0)}=F^{(W)}$, $F^{(+1)}=F^{(A)}$ on $\mathbb{C}$ that arise from $G=\hat{\rho}$, where $\hat{\rho}$ is a specified (single-mode) density operator that is an element of $\mathcal{W}$, or an extension of $\mathcal{W}$ that includes formal power series in $a,{a^{\dagger}}$ [2, 21, 22]. These are called the $Q$, $W$, and $P$ quasiprobability densities, and are attributed respectively to Husimi, Wigner, and Glauber–Sudarshan. But in older works on phase-space methods in quantum optics, operators other than density operators and observables were frequently “classicized,” and it was in that general context that $s$-ordering was introduced. (See [23, 24, 25], and more recently [26]. [^1] There are two conventions on the sign of $s$. In the original paper of Cahill–Glauber [1] and a few more recent ones [26, 27, 28], s=+1,1s=+1,-1 correspond to normal and anti-normal ordering; e.g., for a density operator, to $Q$ and $P$. But usually, they correspond to anti-normal and normal ordering (e.g., to $P$ and $Q$). See, e.g., [2, Chapter 12], [6, 29], and [3, Chapter 3]. The latter convention has been adopted here, so the sign of $s$ in eq. (1.2), taken from [1], has been changed. Note also that in some works (e.g., [21] and [24, Table V]), the meanings of the so-called reciprocal orderings $F^{(N)}$ and $F^{(A)}$, related by negation of $s$, are interchanged.)
@@
Differentiating the exponential operator $D(\beta,\beta^{*};s)={\rm e}^{(s^{\prime}-s)|\beta|^{2}/2}D(\beta,\beta^{*};s^{\prime})$ with respect to $\beta,\beta^{*}$, in a procedure that can be traced to Wilcox [30], and using (2.7), yields the $s$-ordering conversion formula
@@
$\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s}=\sum_{k=0}^{\min(m,n)}k!\,\binom{n}{k}\binom{m}{k}\left(\frac{s-s^{\prime}}{2}\right)^{k}\mathopen{:}\,\mathrel{{\alpha^{*}}^{n-k}\alpha^{m-k}}\,\mathclose{:}_{s^{\prime}}.$ (2.13)
The $n=m=1$ case of (2.13) is the contraction
$\mathopen{:}\,\mathrel{\alpha^{*}\alpha}\,\mathclose{:}_{s}-\mathopen{:}\,\mathrel{\alpha^{*}\alpha}\,\mathclose{:}_{s^{\prime}}=\frac{s-s^{\prime}}{2},$ (2.14)
@@
from which the general formula (2.13) can be derived, much as Wick’s theorem is derived [28]. As another example, the $s=\penalty 10000+1$, $s^{\prime}=-1$ case of (2.13) is
$a^{m}{a^{\dagger}}^{n}=\sum_{k=0}^{\min(m,n)}k!\,\binom{n}{k}\binom{m}{k}{a^{\dagger}}^{n-k}a^{m-k},$ (2.15)
@@
an identity in $\mathcal{W}$ that converts an anti-normal ordering ($s=\penalty 10000+1$) to a normal one ($s=\penalty 10000-1$). Thus, one can convert between the coefficients of an anti-normal and a normal series, $\sum_{m,n}f_{m,n}^{(A)}a^{m}{a^{\dagger}}^{n}$ and $\sum_{m,n}f_{n,m}^{(N)}{a^{\dagger}}^{n}a^{m}$ [31].